OpenStudy (anonymous):

what is the equation in the point-slope from the line passing through (0 2) and (1 5) a) y-5=3(x-1) b) y-2=-3(x-0) c) y+2=3(x+1) d) y+5=-3(x+1)

OpenStudy (anonymous):

\[\frac {y-5}{5-2}=\frac {x-1}{1-0}\]\[\frac {y-5}{3}=x-1\]y-5=3(x-1)

OpenStudy (anonymous):

and the other one please ....

OpenStudy (anonymous):

what is the equation in standard form of the line which passes through (1,-3) and has the slope of 2 a) 2x-y=5 b) 2y-y=-5 c)2x+y=5 d) 2x+y=-5